# SAFER-Splat: A Control Barrier Function for Safe Navigation with Online Gaussian Splatting Maps

### v1

@@ VI Experiments
Simulation Results. We compare SAFER-Splat (green) to NeRF-CBF [19] (red), in addition two variants of our method: one which uses a conservative bounding sphere of each ellipsoid (blue) and one that uses the GSplat means (yellow) in Fig. 3. A Mahalanobis distance variant was also tested but proved too ill-behaved to show. NeRF-CBF is trained using Nerfacto [46] on the same dataset as the GSplat. We examine each method in five scenes: Stonehenge (100K Gaussians), Statues (200K), and Flightgate (300K) (from [14]); in addition to Adirondacks (500K), and low-res Flightgate (15K). All controllers are executed for $100$ trajectories. Our method strikes a favorable trade-off compared to other methods.
The computation time of SAFER-Splat and its variants are similar and more than an order of magnitude faster than NeRF-CBF. We also plot timing results without constraint pruning (*), on a point cloud (about 6$\times$ more points than ellipsoids) represented by $\times$, and a combination of both ($+$), demonstrating that our real-time performance is attributed to both constraint pruning and the sparse GSplat representation.
To assess safety, we compute the squared distance of the trajectories to the high-res GSplat 7 (instead of COLMAP due to the better dense reconstruction from the GSplat). NeRF-CBF attains major violations (data points less than 0) compared to our method and its variants. Our method and its variants can become unsafe due to dynamics noise, but infrequently. We expect that in sparse reconstructions with large, elongated Gaussians, e.g., in the early stages of GSplat mapping, our method strikes a promising tradeoff between the conservativeness of the ball variant and the high potential for collision of the means-only variant. We see this behavior beginning to appear in low-res Flightgate (15K Gaussians).
We utilize the objective of 2 (labeled Control Difference) and progress to the goal ($\min_{t}||\mathbf{p}(t)-\mathbf{p}_{0}||/||\mathbf{p}_{f}-\mathbf{p}_{0}||$) as proxies for conservativeness. We do not expect the progress to be 1 because all methods are reactive and utilize a simple PD controller to navigate to the goal. We observe that NeRF-CBF is sporadic in its performance, with it failing completely on Stonehenge and yielding only a few valid trajectories in Statues and Flightgate. Again, our method and the point variant achieve similar results, which are slightly less conservative than the ball variant.
@@
Hardware Platform. In our hardware experiments, we use a 220 mm Starling 2 drone. The platform maintains a pose estimate using visual inertial odometry (VIO) and can track position, velocity, and acceleration commands with a low-level PX4 controller. The camera images, point clouds, and VIO poses are sent through ROS2 [47] to an off-board desktop computer with RTX 4090 GPU via 5 GHz WiFi. The high-level drone acceleration commands come from a teleoperator and are relayed to the off-board computer which transmits CBF corrected commands back to the drone’s low-level controller. SAFER-Splat’s ROS nodes are made available in our codebase. Note that the scenes used in the real-world experiments contain visual fiducials (ArUco) and motion capture cameras, but these are not used by our VIO or GSplat mapping modules for pose estimation.
@@
Figure 5: Top row: Overlayed color and depth renders from the GSplats produced using SAFER-Splat in the real-world experiments. Bottom row: Overlayed corresponding keyframe color image and the sparse ToF PC.
@@
Hardware Experiments. In three real-world experiments, SAFER-Splat successfully enabled safe teleoperation even when the operator repeatedly attempted to force a collision by driving the drone towards obstacles. Fig. 4 shows our sample trajectories from each scene along with the full plots of our safety value versus time, demonstrating that the safety layer avoids collision ($h_{min}>0$) and only turns on close to the boundary (i.e., minimally invasive). Our method performed online optimization on the parameters of between 150,000 and 300,000 Gaussians in each scene while simultaneously computing safety filtered actions at 15Hz. We constructed our GSplat maps entirely while the drone was flying, and our CBF enabled after a brief warm-up period (5-30 seconds of flight). During flights in Environments 1 and 3, there were brief violations of the CBF, $h_{min}<0$, in all cases these violations were on the order of 1-4mm and due to minor failures of the drone’s low-level controller in accurately tracking the safety layer’s outputs. Likely these tracking failures were caused by rotor induced aerodynamic effects when close to obstacles, and in no cases did they result in loss of control of the drone. Full recordings of each of these real-world flights are available on our project’s website. Fig. 5 showcases the high visual accuracy of each environment’s GSplat when trained via SplatBridge.
VII Conclusion
@@
We propose a real-time 3D reconstruction and safe robot action filtering pipeline called SAFER-Splat for GSplats, which are photo-realistic, efficient to optimize, and interpretable. The proposed safety layer utilizes a CBF formulated for ellipsoidal robot geometries, and leverages the favorable properties of the GSplat representation for scalable and rigorous safety guarantees. SAFER-Splat is generalizable to arbitrary high-level planners, minimally-invasive, and well-behaved, resulting in smooth robot motion. The proposed safety layer is coupled with an online GSplat mapping module to enable safe navigation in unknown environments. In simulations, we show that SAFER-Splat out-performs comparable methods based on NeRFs, and in real-world experiments, we demonstrate that SAFER-Splat can simultaneously optimize a 3D reconstruction and perform as a safe action filter for a teleoperated drone.
VIII Limitations and Future Work
@@
SAFER-Splat assumes that the underlying GSplat map is relatively accurate with respect to the real-world, which does not hold in some situations, especially with dynamic objects. Future work will seek to improve SAFER-Splat in these cases by accounting for dynamic objects in an uncertainty-aware optimization framework. Other future work will aim to improve the robustness of SplatBridge to inaccurate estimated camera poses, which could affect the accuracy of the GSplat.
We also aim in future work to extend SAFER-Splat to semantic mapping and semantic-aware safety. Furthermore, the generalizability of SAFER-Splat to arbitrary high-level planners enables our method to act as a safety filter for, e.g., AI-based Vision-Language-Action (VLA) models.

### v2

@@ VI Experiments
Simulation Results. We compare SAFER-Splat (green) to NeRF-CBF [19] (red), in addition two variants of our method: one which uses a conservative bounding sphere of each ellipsoid (blue) and one that uses the GSplat means (yellow) in Fig. 3. A Mahalanobis distance variant was also tested but proved too ill-behaved to show. NeRF-CBF is trained using Nerfacto [46] on the same dataset as the GSplat. We examine each method in five scenes: Stonehenge (100K Gaussians), Statues (200K), and Flightgate (300K), Adirondacks (500K), and low-res Flightgate (15K). All controllers are executed for $100$ trajectories. Our method strikes a favorable trade-off compared to other methods.
The computation time of SAFER-Splat and its variants are similar and more than an order of magnitude faster than NeRF-CBF. We also plot timing results without constraint pruning ($\star$), on a point cloud (about 6$\times$ more points than ellipsoids) represented by $\times$, and a combination of both ($+$), demonstrating that our real-time performance is attributed to both constraint pruning and the sparse GSplat representation.
To assess safety, we compute the squared distance of the trajectories to the high-res GSplat (7) (instead of COLMAP due to the better dense reconstruction from the GSplat). NeRF-CBF attains major violations (data points less than 0) compared to our method and its variants. Our method and its variants can become unsafe due to dynamics noise, but infrequently. We expect that in sparse reconstructions with large, elongated Gaussians, e.g., in the early stages of GSplat mapping, our method strikes a promising tradeoff between the conservativeness of the ball variant and the high potential for collision of the means-only variant. We see this behavior beginning to appear in low-res Flightgate (15K Gaussians).
We utilize the objective of 2 (Control Difference) and progress to the goal ($\min_{t}||\mathbf{p}(t)-\mathbf{p}_{0}||/||\mathbf{p}_{f}-\mathbf{p}_{0}||$) as proxies for conservativeness. We do not expect the progress to be 1 because all methods are reactive and utilize a simple PD controller ($\bar{u}$) to navigate to the goal. We observe that NeRF-CBF is sporadic in its performance, failing completely on Stonehenge and yielding only a few valid trajectories in Statues and Flightgate. Again, our method and the point variant achieve similar results, which are slightly less conservative than the ball variant.
Figure 4: Top row: Overlaid color and depth renders from the GSplats produced using SAFER-Splat in the real-world experiments. Bottom row: Overlaid corresponding keyframe color image and sparse point cloud.
@@
Hardware Platform. Our experiments are executed on a 220 mm Starling 2 drone. The platform maintains a pose estimate using visual inertial odometry and tracks position, velocity, and acceleration commands with a low-level PX4 controller. Images, point clouds, and VIO poses are sent through ROS2 [47] to an off-board desktop computer with RTX 4090 GPU via 5 GHz WiFi. The high-level drone acceleration commands come from a teleoperator and are relayed to the off-board computer which transmits CBF corrected commands back to the drone’s low-level controller. SAFER-Splat’s ROS nodes are made available in our codebase. Note that the scenes used in the real-world experiments contain visual fiducials and motion capture cameras, but these are not used by our VIO or GSplat mapping modules for pose estimation.
@@
Hardware Experiments. In three real-world experiments, SAFER-Splat successfully enabled safe teleoperation even when the operator repeatedly attempted to force a collision by driving the drone towards obstacles. Fig. 5 shows our sample trajectories from each scene along with the full plots of our safety value versus time, demonstrating that the safety layer avoids collision ($h_{min}>0$) and only turns on close to the boundary (i.e., minimally invasive). Our method performed online optimization on the parameters of between 150,000 and 300,000 Gaussians in each scene while simultaneously computing safety filtered actions at 15Hz. We constructed our GSplat maps entirely while the drone was flying, and enabled our CBF after a brief warm-up period (5-30 seconds of flight). During flights in Environments 1 and 3, there were brief violations of the CBF, $h_{min}<0$, in all cases these violations were on the order of 1-4mm and due to minor failures of the drone’s low-level controller in accurately tracking the safety layer’s outputs. Likely these tracking failures were caused by rotor induced aerodynamic effects when close to obstacles, and in no cases did they result in loss of control of the drone. Full recordings of each of these real-world flights are available on our project’s website. Fig. 4 showcases the high visual accuracy of each environment’s GSplat when trained via SplatBridge.
VII Conclusion, Limitations, and Future Work
@@
We propose a real-time 3D reconstruction and safety filter pipeline using GSplats, which are photo-realistic, efficient to optimize, and interpretable. The proposed safety layer utilizes a CBF formulated for ellipsoidal robot geometries, and leverages the favorable properties of the GSplat representation for scalable and rigorous safety guarantees. SAFER-Splat is generalizable to arbitrary high-level planners, minimally-invasive, and well-behaved, resulting in smooth robot motion. The proposed safety layer is coupled with an online GSplat mapping module to enable safe navigation in unknown environments. In simulations, we show that SAFER-Splat out-performs comparable methods based on NeRFs, and in real-world experiments, we demonstrate that SAFER-Splat can simultaneously optimize a 3D reconstruction and perform safety filtering for a teleoperated drone.
@@
SAFER-Splat assumes that the underlying GSplat map is relatively accurate with respect to the real-world, which does not hold in some situations, especially with dynamic objects. Future work will seek to improve SAFER-Splat in these cases by accounting for dynamic objects in an uncertainty-aware optimization framework. Second, we hope to expand SAFER-Splat beyond double-integrator systems, e.g., contact-rich dynamical systems. Next, SAFER-Splat can be deployed on lower-end hardware as GSplat maps become more lightweight [48]. Furthermore, SplatBridge can be made more robust to inaccurate camera pose estimates, which affect the accuracy of the GSplat. Finally, we seek to extend SAFER-Splat to semantic multi-robot mapping [49, 50], semantic-aware safety, and simultaneous deployment with high-level planners like Vision-Language-Action models.